OpenStudy (anonymous):

if the point p(-3/5, y) lies on the unit circle and P is the second quadrant what does y equal?

OpenStudy (amistre64):

y = sin(t) = sqrt(1-cos(t)^2)

OpenStudy (amistre64):

where cos(t) = -3/5 of course :)

OpenStudy (amistre64):

ima bet its 4/5 just to enigmatic and all

OpenStudy (anonymous):

i think so too the 4/5 answer

OpenStudy (amistre64):

:) the 3-4-5 triangle is a favorite